Construct an orthonormal basis for the subspace of R3 spanned by the vectors (1,-1,1)

and (2,0,4)


1
Expert's answer
2021-08-17T17:50:48-0400

Let us find "a\in\R" such that the vectors "\overline{u}=(1,-1,1)" and "\overline{v}=(1,-1,1)+a(2,0,4)=(1+2a,-1,1+4a)" are orthogonal. Then the dot product of these vectors is equal to 0, that is "1+2a+1+1+4a=0." It follows that "6a=-3," and hence "a=-\frac{1}{2}."

Then "\overline{v}=(1+2(-\frac{1}{2}),-1,1+4(-\frac{1}{2}))=(0,-1,-1)." Since "|\overline{u}|=\sqrt{1+1+1}=\sqrt{3}" and "|\overline{u}|=\sqrt{0+1+1}=\sqrt{2}," we conclude that the vectors "\frac{\overline{u}}{|\overline{u}|}=(\frac{1}{\sqrt{3}},-\frac{1}{\sqrt{3}},\frac{1}{\sqrt{3}})" and "\frac{\overline{v}}{|\overline{v}|}=(0,-\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}})" form an orthonormal basis for the subspace of "\R^3" spanned by the vectors "(1,-1,1)" and "(2,0,4)."